# Senquan Wang

### 12 Mar 2024 v1

@@ Abstract
Blood Glucose (BG) control involves keeping an individual’s BG within a healthy range through extracorporeal insulin injections is an important task for people with type 1 diabetes. However, traditional patient self-management is cumbersome and risky. Recent research has been devoted to exploring individualized and automated BG control approaches, among which Deep Reinforcement Learning (DRL) shows potential as an emerging approach. In this paper, we use an exponential decay model of drug concentration to convert the formalization of the BG control problem, which takes into account the delay and prolongedness of drug effects, from a PAE-POMDP (Prolonged Action Effect-Partially Observable Markov Decision Process) to a MDP, and we propose a novel multi-step DRL-based algorithm to solve the problem. The Prioritized Experience Replay (PER) sampling method is also used in it. Compared to single-step bootstrapped updates, multi-step learning is more efficient and reduces the influence from biasing targets. Our proposed method converges faster and achieves higher cumulative rewards compared to the benchmark in the same training environment, and improves the time-in-range (TIR), the percentage of time the patient’s BG is within the target range, in the evaluation phase. Our work validates the effectiveness of multi-step reinforcement learning in BG control, which may help to explore the optimal glycemic control measure and improve the survival of diabetic patients.
@@ 1 Introduction
Diabetes profoundly affects human life and health, regardless of country, age or gender, and is one of the leading causes of death and disability worldwide [1]. From 1990 to 2021, the age-standardized prevalence of diabetes increased by 90.5 % globally, with increases of more than 100 % in several regions, and it is projected that by 2050 there will be 1.31 billion people with diabetes worldwide [1]. Furthermore, people with diabetes have more than twice the normal risk of early death, resulting in an estimated 150-500 million deaths around the world each year, while generating approximately 12% of health expenditure (\$966 billion) [2, 3]. The rising prevalence and serious health and economic hazards have attracted the attention of scientists around the globe, and as a result, the number of studies on diabetes is increasing.
@@
The pancreas of a diabetic does not produce or produces very little insulin or the insulin produced is not used efficiently leading to high BG and a variety of life-threatening complications such as cardiovascular disease, nerve damage, kidney damage, lower limb amputations, and eye disease leading to decreased vision and even blindness [3]. BG control is their basic treatment, as well as the basis for the prevention and treatment of diabetic complications. Patients mainly control the stability of BG by injecting insulin. However, this traditional self-management is usually cumbersome and challenging, as it requires patients to measure their BG levels several times a day, while they suffer from many of the aforementioned complications [2].
Another potential approach to glucose control is using Reinforcement Learning (RL) algorithms, a branch of machine learning suitable for solving sequential decision problems [4]. RL makes decisions based on observed states, requiring neither a detailed description of the environment nor a large amount of labeled data for training as supervised learning does, and can learn directly on real data or interact with dynamic systems represented by mathematical models to automatically derive personalized dosing regimens for patients with life-threatening diseases [4, 5]. Furthermore, the lagged effects of actions in RL and the long-term presence of rewards are precisely adapted to the time lag in subcutaneous glucose measurements and insulin injections. [2, 4]. The prolongedness of drug effects is also a factor that requires special consideration, although only a few studies have explored this, as ignoring delayed and prolonged effects of drugs can lead to toxicity associated with overdosage [6].
Currently, the combination of deep learning with outstanding representational capabilities and RL with strong decision-making capabilities as DRL has more significant potential to extend RL to previously intractable problems [7]. One of the most exemplary representatives of DRL is the Deep Q-Network (DQN)[8], which can learn successful policies directly from high-dimensional inputs using end-to-end RL and is the first artificial agent capable of learning to excel in a variety of challenging tasks. So far, many extensions have been proposed to improve the learning speed or stability of DQN. For example, Double DQN [9] (DDQN) to reduce overestimation of DQN, PER [10] to enable DQN to learn more efficiently, Dueling network architecture [11] to allow DQN to generalize learning across actions, multi-step Q-learning [12, 13], Distributional Q-learning [14], and Noisy DQN [15], etc. The integration of all these components into a single integration agent is called Rainbow [16].
In this article, we use the DRL algorithm DQN and its variants PER and multi-step learning to achieve individualized BG control in virtual type 1 diabetes patients. The delayed and prolonged effect of the drug was also considered and the problem of BG control was converted from a PAE-POMDP to a MDP through a drug concentration decay model [6]. Based on the modified Simglucose model of Basu et al. our proposed improved strategy performs better than the benchmark in both the training and evaluation phases. The remainder of this paper is organized as follows: section 2 enumerates the relevant work on RL in healthcare, especially for BG control. Section 3 presents background knowledge of the decision-making process and multiple variants of DQN. Section 4 provides the technical approach. Section 5 shows the results of training and evaluation. Section 6 states the conclusions.
2 Related Work
In recent years, RL in healthcare has seen an increase in research and successful applications in several areas of disease, such as sepsis [17, 18], anemia [19], schizophrenia [20], and intravenous heparin infusion [21] and pain management [22] in critically ill patients in the intensive care unit. In particular, closed-loop BG control in patients with type 1 diabetes mellitus, a long-standing challenge in medicine, has inspired RL researchers to explore it for a long time [6]. Relevant studies from 1990 to 2019 can be seen in the review by Tejedor et al. [2]. They did an exhaustive literature search and evaluation, ultimately filtering out 29 papers that focused strictly on glucose control and used RL methods. The review reports that actor-critic learning, the most popular RL algorithm, was used in 11 papers, while only 1 used the DQN algorithm [23]. This suggests that there is still much room to explore the application of DQN algorithms in glucose control. Later, Fox et al. used the Soft Actor-Critic (SAC) algorithm to learn a strategy for glucose control, and they framed the closed-loop glucose control problem as a Partially Observable Markov Decision Process (POMDP) by taking into account the inherent noise of the continuous glucose monitor (CGM) sensors and the unobserved meal assumption [24]. Zhu et al. also used the SAC algorithm and introduced food information through the Gate Recurrent Unit (GRU) [25]. Basu et al. state that the prolongedness of action effects is the cause of the partial observability of the state and convert PAE-POMDPs into MDPs by exponential decay of the effective action [6]. Our work focuses on improving the performance of BG control based on the work of Basu et al. using PER and multi-step learning strategies.
@@ 2 Related Work
Simple Q-Networks have several shortcomings [17]. The use of many DQN extensions can be seen in drug dosing problems, and they can often obtain more promising results than the classical DQN algorithm. For example, Raghu et al. combined the DDQN, Dueling DQN framework, and PER methods to produce the Dueling Double-Deep Q-Network (D3QN) to learn optimal sepsis treatment strategies [17]. Lopez-Martinez et al. also used D3QN to learn optimal opioid administration strategies for the management of critical care pain [22]. Likewise, the use of DQN extensions can be found in studies of glucose control problems, e.g., Zhu et al. conducted a study of monohormonal (insulin) and dual-hormonal (insulin and glucagon) administration for glucose control using DDQN as a dosing strategy [26]. Several deep q-learning algorithms were thoroughly tested and compared by Tejedor et al [27]. However, they did not take into account the prolongedness of the action effects and the multi-step learning approach.

### 15 Mar 2024 v2

@@ Abstract
Blood Glucose (BG) control involves keeping an individual’s BG within a healthy range through extracorporeal insulin injections is an important task for people with type 1 diabetes. However, traditional patient self-management is cumbersome and risky. Recent research has been devoted to exploring individualized and automated BG control approaches, among which Deep Reinforcement Learning (DRL) shows potential as an emerging approach. In this paper, we use an exponential decay model of drug concentration to convert the formalization of the BG control problem, which takes into account the delay and prolongedness of drug effects, from a PAE-POMDP (Prolonged Action Effect-Partially Observable Markov Decision Process) to a MDP, and we propose a novel multi-step DRL-based algorithm to solve the problem. The Prioritized Experience Replay (PER) sampling method is also used. Compared to single-step bootstrapped updates, multi-step learning is more efficient and reduces the influence of biasing targets. Our proposed method converges faster and achieves higher cumulative rewards compared to the benchmark in the same training environment, and improves the time-in-range (TIR), the percentage of time the patient’s BG is within the target range, in the evaluation phase. Our work validates the effectiveness of multi-step DRL in BG control, which may help explore the optimal glycemic control measure and improve the survival of diabetic patients.
@@ 1 Introduction
Diabetes profoundly affects human life and health, regardless of country, age, or gender, and is one of the leading causes of death and disability worldwide [1]. From 1990 to 2021, the age-standardized prevalence of diabetes increased by 90.5 % globally, with increases of more than 100 % in several regions, and it is projected that by 2050, there will be 1.31 billion people with diabetes worldwide [1]. Furthermore, people with diabetes have more than twice the normal risk of early death, resulting in an estimated 150-500 million deaths around the world each year, while generating approximately 12% of health expenditure (\$966 billion) [2, 3]. The rising prevalence and serious health and economic hazards have attracted the attention of scientists around the globe, and as a result, the number of studies on diabetes is increasing.
@@
The pancreas of a diabetic does not produce or produces very little insulin, or the insulin produced is not used efficiently, leading to high BG and a variety of life-threatening complications such as cardiovascular disease, nerve damage, kidney damage, lower limb amputations, and eye disease leading to decreased vision and even blindness [3]. BG control is their basic treatment, as well as the basis for preventing and treating diabetic complications. Patients mainly maintain the stability of BG by injecting insulin. However, this traditional self-management is usually cumbersome and challenging, as it requires patients to measure their BG levels several times a day, while they suffer from many of the aforementioned complications [2].
To reduce this burden on diabetic patients and better help them with BG control, the Artificial Pancreas (AP) system has been proposed as an effective and safe new approach to BG management and has become the focus of diabetes research [4]. The AP consists mainly of three parts: a Continuous Glucose Monitoring (CGM) system, an insulin pump, and BG control algorithm. The closed-loop control algorithm is the core of the AP and continues to evolve. Previously, three main types of algorithms, namely the Proportional Integral Derivative (PID) control algorithm, Model Predictive Control (MPC) algorithm, and Fuzzy Logical (FL) control algorithm have been used in closed-loop BG control research. The PID controller is reactive and is unable to react fast enough to meals [5, 6]. MPC builds systems based on assumptions that may ignore some unknown variables resulting in reduced accuracy [7]. The FL approach to modeling expert regulation of insulin delivery based on approximation rules may not provide customized BG control strategies for patients [8]. Overall, these traditional reactive controllers based on instantaneous glucose changes have not completely eliminated the effects of the delay inherent in the AP system predisposing to postprandial hyperglycemia, and the variability in BG concentrations due to dietary intake, exercise, sleep, and stress has not yet been effectively modeled, making it difficult to achieve patient-individualized precise glycemic control [9].
In recent years, the Reinforcement Learning (RL) algorithm has been recognized as one of the most promising approaches capable of replacing the traditional paradigm of controlling insulin infusion, and it can be used to design a fully closed-loop controller that provides a truly personalized BG control scheme based entirely on the patient’s own data [2]. The BG control problem, which focuses on deciding the insulin dose to be injected into the patient at each moment, is a time-series decision-making problem, and RL algorithms are classical approaches for solving such sequence decision-making problems. RL makes decisions based on observed states, requiring neither a detailed description of the environment nor a large amount of labeled data for training as supervised learning does, and can learn directly on real data or interact with dynamic systems represented by mathematical models to automatically derive personalized dosing regimens for patients with life-threatening diseases [7, 10]. The review by Tejedor et al. evaluates the use of RL methods to design algorithms for glycemic control in diabetic patients from 1990 to 2019 and predicts that RL algorithms will be used more frequently for BG control in the coming years [2]. Fox et al. implemented automated BG control using three deep Q-networks that differ in architecture and state representation [11]. Later, Fox et al. constructed the closed-loop BG control problem as a Partially Observable Markov Decision Process (POMDP), then used the Soft Actor-Critic (SAC) algorithm to learn a strategy for BG control, taking into account the noise of the CGM sensors and unobserved meal assumptions, and emphasized the flexibility of the RL methodology by experimentally comparing it with non-reinforcement learning algorithms [5]. Nordhaug Myhre et al. used a policy gradient RL approach for automated BG control, achieving performance that outperforms even state-of-the-art MPC [12]. Zhu et al. also used the SAC algorithm and introduced food information through the Gate Recurrent Unit (GRU) [13]. Di Felice et al. compared the performance of two algorithms, DDPG and SAC, on a BG control task, with the latter achieving superior performance [14]. Basu et al. considered the delay and prolongedness of drug effects and made improvements based on the DQN algorithm, finally proposing Effective-DQN algorithm to achieve individualized control of BG [15].
@@
The combination of RL and deep learning with outstanding representational capabilities is being brought to its greater potential to scale to previously intractable problems [16]. One of the most exemplary representatives of DRL is the Deep Q-Network (DQN)[17], which can learn successful policies directly from high-dimensional inputs using end-to-end RL and is the first artificial agent capable of learning to excel in a variety of challenging tasks. To date, many extensions have been proposed to improve the learning speed or stability of DQN. For example, Double DQN [18] (DDQN), PER [19], Dueling network architecture [20], multi-step Q-learning [21, 22], Distributional Q-learning [23], and Noisy DQN [24], etc., all of the above components are integrated into a single integration agent is called Rainbow [25]. The use of many DQN extensions can be seen in drug dosing problems, and they can often obtain more promising results than the classical DQN algorithm. For example, Raghu et al. combined the DDQN, Dueling DQN framework, and PER methods to produce the Dueling Double-Deep Q-Network (D3QN) to learn optimal sepsis treatment strategies [26]. Lopez-Martinez et al. also used D3QN to learn optimal opioid administration strategies for the management of critical care pain [27]. Likewise, the use of DQN extensions can be found in studies of glucose control problems, e.g., Zhu et al. conducted a study of monohormonal (insulin) and dual-hormonal (insulin and glucagon) administration for glucose control using DDQN as a dosing strategy [28]. Emerson et al. explored three approaches, Batch Constrained Deep Q-learning (BCQ), Conservative Q-learning (CQL) and Twin Delayed DDPG with Behavioral Cloning (TD3-BC), for the management of BG, demonstrating the potential of offline RL to create safe and sample-efficient glycemic control strategies in patients with type 1 diabetes. Several deep q-learning algorithms were thoroughly tested and compared by Tejedor et al [9]. However, they did not consider the prolongedness of the action effects and the multi-step learning approach.
Most previous articles using RL for BG control have used single-step reward bootstrapping; however, the target for single-step updating may be biased and lead to inefficient learning. Multi-step learning can reduce this bias and learn efficiently, yielding better performance. Therefore, in this article, we use a DQN algorithm extended with a multi-step learning approach to implement personalized BG control for virtual type 1 diabetic patients and incorporate the PER sampling method to improve learning efficiency. We also consider the delayed and prolonged drug effects and use a drug concentration decay model to convert the BG control problem from a PAE-POMDP to a MDP. In the modified Simglucose environment, our proposed improved strategy outperforms the benchmark in both the training and evaluation phases. The remainder of this paper is organized as follows: section 2 presents background knowledge of the decision process and the algorithmic mechanics of DQN and its several variants. Section 3 provides the technical approach. Section 4 shows and analyzes the results of training and evaluation. Finally, conclusions and promising future directions are offered in section 5.